### v1

# Formal Definition and Implementation of Reproducibility Tenets for Computational Workflows

Abstract
Computational workflow management systems power contemporary data-intensive sciences. The slowly resolving reproducibility crisis presents both a sobering warning and an opportunity to iterate on what science and data processing entails. The Square Kilometre Array (SKA), the world’s largest radio telescope, is among the most extensive scientific projects underway and presents grand scientific collaboration and data-processing challenges. This work presents a scale and system-agnostic computational workflow model and extends five well-known reproducibility tenets into seven defined for our workflow model. Subsequent implementation of these definitions, powered by blockchain primitives, into the Data Activated Flow Graph Engine (DALiuGE), a workflow management system for the SKA, demonstrates the possibility of facilitating automatic formal verification of scientific quality in amortized constant time. We validate our approach with a simple yet representative astronomical processing task; filtering a noisy signal with a lowpass filter with both CPU and GPU methods. Our framework illuminates otherwise obscure scientific discrepancies and similarities between principally identical workflow executions.
keywords:
Scientific workflows, Scientific reproducibility, Workflow management systems
journal: Elsevier
[…]
[icrar]organization=International Centre for Radio Astronomy Research, University of Western Australia, addressline=7 Fairway, city=Perth, postcode=6009, state=WA, country=Australia
1 Introduction
Reuse is the ultimate goal of science; by building trust in others’ observations, theories, models, and experiences, humanity explores the universe. The ability to reproduce elements of scientific works increases our trust in their findings. The reproducibility crisis [1] heightened scrutiny and discussion around what highly trusted modern science entails. The SKA is an already storied behemoth of academic collaboration that demands paying attention to scientific reproducibility. Radio astronomy relies heavily on computational methods but is fundamentally an experimental, observational science, and hence, it presents a curious case for scientific reproducibility. Seeking a balance between the world of formal software engineering and the exploratory nature of scientific investigation has proved challenging to achieve [2, 3]. The SKA, representing an order of magnitude increase in data-processing requirements on current observatories [4], benefits from a focus on building scientific reproducibility into its instrumentation, providing an opportunity to improve what reproducibility practically means for astronomy and other computationally powered sciences. The literature widely debates the precise meaning of scientific reproducibility and replicability [5, 6] and, concerning computing, focuses on ‘forwards facing’ approaches to guarantee the future reproducibility of current computations. While practically valuable, the general trend toward adhering to specific technologies and conventions is potentially ruinous from a philosophical standpoint; how can we know that our efforts to make scientific works reproducible succeed? This work presents and validates a scale and system agnostic framework for ‘backwards facing’ reproducibility tests on computational workflow executions, facilitating efficient verification of scientific reproducibility for astronomical workflows. Specifically, we make the following contributions:
  1. 1.
    Present a scale-agnostic computational workflow model suitable for computational reproducibility discussions.
  2. 2.
    Define seven reproducibility tests that we term tenets, based on known literature.
  3. 3.
    Describe a hash-graph-based workflow signature method, termed BlockDAGs, based on the scale-agnostic model allowing for constant time assessment of our reproducibility tenets.
  4. 4.
    Implement this signature mechanism in the DALiuGE [7] workflow management system.
  5. 5.
    Demonstrate the efficacy of our model, workflow signature method, and reproducibility tenets with a demonstrative lowpass filter workflow. Our method uncovers obscure differences between in-principal identical workflows.
2 Related Work
Reproducibility is fundamental to science, and computational reproducibility in support of quality science is a long-standing goal in the field, but one where the very concept is confused. Computational workflow managers are among the most powerful tools available in this quest, which, alongside contemporary software engineering techniques and provenance capture, are moving towards a world of fully rerunnable, repeatable, reproducible, and replicable research. Claerbout [8] coined the concept of computational reproducibility in scientific computing, and Peng [9] elaborates on what is needed to achieve computational reproducibility in science. Drummond [10] separates the concepts of reproducibility and applicability. Barba [5] reviews the myriad of terminologies used to describe computational reproducibility as Benureau and Rougier [11] provide a framework separation rerunning, repeating, reproducing, and replicating computational works. Central to these discussions is a separation between bitwise precise replication of exact results and conceptual reproductions of an existing work independent of precise implementation. Our definitions of these reoproducibility tenets unify these two views. Current suggested best practice in pursuit of scientific reproducibility range simple rules [12], adapting modern tooling from industry including notebook environments [13, 14, 15], sharing code [16] and publishing openly [17, 18]. Related to the signature approach we present in this work, Bellini investigates the feasibility of building a DOI-like dataset identifier leveraging public blockchains [19].
Computational reproducibility is a driving motivation in creating scientific workflow management systems. In the world of computational workflow management systems, reproducibility often relates to storing workflow descriptions as scientific artifacts themselves [20, 21], according to specific standards such as FAIR [22] for later investigation and use. da Silva et al. [23] cite dynamic provenance capture as necessary for workflows to scale larger and tackle more challenging scientific problems. Gaignard et al. [24] leverage semantic web technologies to produce valuable workflow provenance information, later expanding this concept into a workable ontology [25].
Highly scalable and trustable computing is critical to conquering our largest scientific endeavors [26], and as we move towards reproducible computational workflow management, we must also ensure our efforts function as expected; a key motivation behind our signature-based approach to workflow reproducibility.
3 Computational Workflows
Computational workflows (herein, ‘workflows’) are structured assemblies of computational tasks and data stores [27] and underpin some of the world’s most extensive science projects, including the SKA [4]. Workflows separate domain-specific software knowledge from scientific methodology, making increasingly complex computations possible for increasingly more scientists, and have seen broad uptake in the past three decades [28]. Many workflow management systems provide several flavors of workflows for specific science domains [29], and thus, any attempt to address workflow reproducibility should be tool-agnostic. DALiuGE [7], [30] is a workflow management system developed to execute workflows at the scale the SKA requires. The unique data-driven design, decentralized execution, and a careful share-nothing architecture enable this. Moreover, DALiuGE strictly separates logical workflow design from physical workflow execution and thus allows astronomers to design workflows without intimate component knowledge and, more importantly, facilitates the broad reuse of logical workflows. Derived from the DALiuGE design, our scale and system-agnostic generalized workflow model alternates data artifacts with computational tasks and separates logical elements from physically executed elements. This separation provides more flexibility in reasoning about reproducing a workflow execution.
3.1 Scale and System Agnostic Workflow Model
Figure 1 presents the generalized Unified Modeling Language (UML) workflow model to make our reasoning about testing workflow executions concrete and applicable to arbitrary workflow management systems.
Figure 1: A UML diagram depicting our arbitrary-scale workflow definition.
Scientific information effectively expresses the information used in making a scientific claim and is derived by running a workflow.
A workflow is a collection of components and data artifacts with an imposed ordered structure.
A logical workflow is a workflow comprised of logical components.
A physical workflow is a workflow comprised of physical components.
A component is an atomic digital entity; it could be a (logical or physical) task or data artifact.
A logical task describes a task characterized by details like programming language, algorithm, and required options and parameters.
A physical task is a single, executable instance of a logical task. A logical task could be realized by one or more physical tasks depending on the degree of parallelism of the workflow.
A logical data artifact is a logical data resource characterized by storage type and required configurations.
A physical data artifact is the actual datastore like a file, distributed file system, or database. A single logical data artifact may encompass many individual physical data artifacts at runtime, dependent on the degree of parallelism of the workflow.
This ordered structure requires components and data artifacts to appear in alternating order; another component cannot precede a component.
A UML definition makes all modeled workflows well-defined software in their own right and permits native workflow implementation independent of using a workflow management system. Moreover, we separate logical and physical workflow components. This separation means we can reason and compare logical workflow designs irrespective of their eventual implementation in software as physical components. Moreover, several physical components can implement a single logical workflow component, as is the case in large-scale workflows. Finally, our model’s requirement to alternate computing tasks and data artifacts ensures that a workflow description captures all interim and transient data artifacts. We note that this model applies only to conventional task-based workflow management systems. While further model formalization, including ontologies for generalized scientific processes, allows accurate descriptions of data provenance [31], further formalization is beyond the scope of this research, whose focus lies on direct software implementation and testing of high-level reproducibility concerns.
3.1.1 Mapping the generalized workflow model to DALiuGE
For this study, we have implemented the generalized ideas described above in DALiuGE since we have complete code control and in-depth knowledge of the code structure. DALiuGE specifies workflows as graphs of components and separates workflows into two main layers, logical and physical, and then sub-divides each of those into template and non-template versions. Workflows are developed as logical templates comprised of exclusively logical components and control structures. Scientists will design a Logical Graph Template (LGT) several months or years before the actual processing and for extensive reuse. LGTs expose configuration options to make them more specific for concrete execution. A Logical Graph (LG) is a LGT with a concrete configuration. LG specification occurs on the order of weeks before execution. Isolating details of storage components only yields a Physical Graph Stencil (PGS), which is defined only for the purpose of provenance storage and is technically a sub-graph of a LG. DALiuGE then unrolls the LG into Physical Graph Template (PGT), resolving control structures and presenting the net computation required. This process is unique for each LG, but the result can vary depending on the translator algorithm used. A physical graph template is a directed acyclic graph (DAG), a common representation for computational workflows [23]. During the scheduling process, DALiuGE partitions a PGT into sub-graphs based on available compute resources, resulting in a Physical Graph (PG). This step is performed just before execution since it maps the physical workflow components to the available resources. Each component of the DAG is a fully specified computational task or data artifact. Finally, the DALiuGE daemon triggers the first physical components in a graph, but after, components communicate with each other in a fully decentralized manner. Components can store runtime-specific information such as log files or data summaries, which together form a Runtime Graph (RG) after graph execution has been completed.
Table 1: DALiuGE to workflow model element mapping.
Workflow Model Element DALiuGE Element
Component Component
Logical Task Logical Component
Physical Task Application Component
Logical Data Artifact Logical Data Component
Physical Data Artifact Data Component
Logical Workflow Logical Graph
Physical Workflow Physical Graph
Table 1 summarizes the mapping of our generic workflow model to DALiuGE components. Adding this layer of abstraction permits multiple workflow system mappings, reproducibility definitions, and testing mechanisms to apply to any mapped system.
4 Reproducible Workflows
Contemporary science is complex; variability in the design and execution of complex data analysis pipelines can be scientifically significant [32]. Most attempts to ensure workflows are reproducible involve analyzing artifact provenance, describing precisely how pieces of data came to exist [31]. We build on this lineage, but instead of designing provenance capture to describe the entire technical process, we formulate reproducibility tests, singling out parts of workflow executions expected to be kept constant between executions. Doing so allows for flexible workflow characterization and does not limit our approach to a specific workflow management system, a critical conceptual requirement for the multi-decade-spanning SKA project and the broader scientific field. Exploiting the structured nature of workflows provides constant time provenance comparison, suitably efficient for deployment at extreme data scales. Our example workflow investigates common difficulties encountered when reproducing computational workflows. A lowpass filter workflow implemented with various methods focuses on a single component and, in the case of Fast Fourier Transform (FFT) implementations, utilizes an operation critical to many astronomical workflows. Our selected example workflows are indicative of common difficulties encountered when reproducing computational workflows.
4.1 Reproducibility Tenets
Melding computing into satisfyingly reproducible science is a well-discussed and unsolved problem [1, 33, 6]. Current efforts range from asserting standards on individual pieces of code [11], enacting well-defined data management protocols [22, 34], containerizing whole workflows [35], measures on behalf of journal publishers [36], and conference organizers [37] and careful provenance production and collection [31]. While certainly practical, we propose that forward-facing measures insufficiently integrate computing into science; they are ultimately empirical and require an inductive step to evaluate their efficacy. Our proposed approach demarcates implementation from abstract computational ideas, data and method reproduction, workflow similarities between workflow management systems, and any combination of these traits. We present seven reproducibility tenets formulated as workflow component invariants, each defining a testable assertion rather than a series of prescribed rules. For each tenet, we provide an unambiguous definition, justification, and a few motivating example scenarios, providing practical motivation behind our definition.
Table 2: Reproducibility tenets summarized as workflow component invariants.
Log. Tasks Phys. Tasks Data Artifacts Scientific Information
Rerun
Repeat
Recompute
Reproduce
Scientific Replication
Computational Replication
Total Replication
We use the popular terminology to rerun, repeat, recompute, reproduce, and replicate computational workflows as these terms capture intuitive concepts around scientific reproducibility [38, 11, 5, 39]. Table 2 summarizes our definitions, and the later section 5.2 describes the precise provenance information the chosen workflow system has to include to satisfy each tenet.
4.1.1 Rerun
A workflow reruns another if they execute the same logical workflow; that is, their logical components and dependencies match.
Rerunning has been suggested to show that used tools are robust; rerunning a computational workflow is analogous to using an identical algorithm independent of its implementation [39]. Asserting consistency only on the logical components in a workflow allows physical component changes. This definition, in turn, allows changes in computing scale, parallelism patterns, hardware, data-encoding, or software implementation while still constituting a rerun. If all physical details are kept identical, a successful rerun builds trust in the underlying tooling used to deploy a workflow, and a failed rerun demonstrates system failure. Rerunning workflows is useful when scaling up from a laptop to a cluster, for instance, or for workflows that, by design, deal with wildly varying data inputs, such as an ingest pipeline.
4.1.2 Repeat
A workflow repeats another if they execute the same logical workflow and a principally identical physical workflow; their logical components and dependencies and physical tasks match. The specificity of a ‘principally identical’ physical workflow allows for a successful recovery from failed physical tasks (due to hardware or operating failure).
Repetition is a relatively well-defined concept, and we generally follow intuition. Repetition builds statistical power behind results [39]. By analogy to laboratory work, repeated scientific computational workflows should match all components (workflow specification, data input, and runtime environment) [38]. Our definition mostly follows existing interpretations but provides a concrete definition. Repeating a scientific computational workflow asserts the reuse of the logical and physical workflows, which, if successful, allows the bundling of data artifacts in concert. For instance, survey analysis runs on many patches of sky to create a full catalog, repeating the same workflow. Repetition has also been known as methods reproducibility [40]. Importantly, we leave the amount of information clarifying physical tasks as an environment and problem-dependent decision but insist that hardware details are not pertinent for repetition.
4.1.3 Recompute
A workflow recomputes another if they execute the same physical workflow; that is, their physical tasks and dependencies match precisely.
Recomputation is a stricter interpretation of computational workflow repetition that requires identical physical tasks and hardware and is not fault-tolerant. Workflow provenance, taken to an albeit extreme conclusion, motivates this definition; the effort required to recompute a given workflow is immense and is likely impossible without total hardware control. Some forms of repetition [11] are similar as they call principally for removing non-determinism. In distributed computational workflows, the distribution itself can cause remarkably complex numerical non-determinism that is painstaking to correct [41], so recomputation is a lofty endeavor. Recomputation can be useful, however, to demonstrate that a new computing facility is stable and to isolate platform-dependent difficulties or errors.
4.1.4 Reproduce
A workflow reproduces another if their scientific information match.
Scientific reproduction is a widely discussed topic, and our definition does not conflict with existing literature. Our definition of reproduction captures the net data flow a computational data-processing workflow consumes and produces. Generally, reproducibility proper refers to reproducing experimental results [42], an extremely difficult task in lab-based science [43]. Some define reproducible codes to be trivially executable and to produce published results [11]. Our definition focuses solely on data in line with other interpretations [39, 5] as opposed to more conceptual definitions based on conclusions [38]. The many interpretations of reproducibility present a challenge with many valid but incompatible definitions [6].
Focussing on data artifacts specifically provides an actionable approach to testing scientific reproducibility. The formal reproduction of other teams’ work and the reproduction of historical but impactful results are healthy for the wider academic community. Our approach is also compatible with reproducibility approaches involving constrained computing environments [44] since we place no constraint on logical or physical task components.
4.1.5 Replicate
Replicability in science is a substantially debated concept with many interpretations [42, 6]. We incorporate replication into our framework as combinations of our previous tenets, creating three different but related standards that all constrain workflow tasks and data artifacts.
A workflow scientifically replicates another if they rerun and reproduce each other.
Our definition asserts that a principally identical computational workflow produces the same terminal data artifacts, in effect verifying that a logical workflow can produce a given set of results. Our interpretations are reminiscent of what is termed conceptual interpretation [42] and of replicable code [11]. Related but alternate views [38] differ in that we do not focus on the conclusions a workflow implies. For example, scientific replication is useful when updating library versions for a particular workflow, using different implementations for the same tasks or moving to a new computing facility.
A workflow computationally replicates another if they recompute and reproduce each other and all data artifacts match.
Our definition asserts that a precisely identical computational workflow produces the same data artifacts at every stage, which is the strictest interpretation of replication similar to the concept of direct replication [42], and is closely related to workflow provenance discussed widely in computational workflow literature [34]. The information required to replicate a workflow execution computationally is immense. It demonstrates that a specific workflow implementation on specific hardware yields a given set of results, building confidence in workflow implementation but not necessarily the underlying science; it would be challenging for an independent team to totally replicate a complex workflow. Recomputation can confirm anomalous or unexpected results before further verification, for instance.
A workflow totally replicates another if they repeat and reproduce each other and all data artifacts match.
Our definition asserts that a practically identical computational workflow produces the same data artifacts, including interim data, and are identical down to software, but not hardware, details. Replication is useful to certify a workflow as a ‘gold-standard’ [39] and is similar to the concept of results reproducibility [40]. For example, in radio astronomy, an integral processing task requires separating unwanted radio frequency interference from astronomical observations, known as RFI flagging. Not only is the correctness of this task important, but the speed of execution also matters significantly, as this task occurs before any science processing can commence. There exist a multitude of possible methods spanning traditional algorithms and machine learning and a performance claim around a new RFI flagging method could be strengthened by replicating the result independently on different hardware. A total replication permits independent verification of results by a third party and balances scientific and computational replication.
5 Methods
Having provided the theoretical background in the previous section, in this section, we outline our practical implementation, which allows us to verify each of the reproducibility tenants. This includes the specific details behind our workflow signature mechanism, the information captured to build these signatures in DALiuGE, and the details of our demonstrative lowpass filter workflow.
5.1 Blockchain-Based Workflow Signature
Building a dynamic workflow signature is central to our reproducibility tests. Moreover, at SKA scale, these signatures need very efficient construction and comparison routines. To this end, we take inspiration from blockchain technologies, borrowing the Merkle tree cryptographic data structure [45]. Information about each workflow model component, dictated by our tenet definitions, is stored in a Merkle tree, the root of which provides a signature for that component. Inserting these signatures into a blockchain mirroring their connectivity in the original workflow creates a block directed acyclic graph (BlockDAG). The root of a Merkle tree constructed from the leaves of this BlockDAG is a signature for the entire workflow. Similar to how blockchains are cryptographically secure, subtle changes to component provenance information result in drastic changes to the final signature, but structuring this signature makes it efficient to identify where these changes occur. Given two workflow BlockDAGs, one can walk through a topological sort of the nodes to find the first differing signature.
We use Kahn’s algorithm [46] to traverse a workflow description in a topological order taking $\mathcal{O}(V+E)$ time where $V$ is the number of components (vertices) in the workflow and $E$ is the number of edges between all components. Construction of each component Merkle tree takes $\mathcal{O}(|V|\,log\,|V|)$ time where $\|V\|$ is the volume of provenance information the component stores. Inserting a component into the BlockDAG takes $\mathcal{O}(d\,log\,d)$ time, where $d$ is the degree of that component. Constructing the entire BlockDAG takes $\mathcal{O}(V(D\,log,D)+E)$ time, where $D$ is the average vertex degree. Construction of the final signature takes $\mathcal{O}(l\,log\,l)$ time, where $l$ is the number of leaves in the workflow graph. When amortized against the time taken to execute a workflow in the first place ($\mathcal{O}(V+E)$), signature construction takes $\mathcal{O}(log\,l)$ time.
5.2 Tenet Data
For each of our tenets, we elaborate on what information specifically is important in the case of DALiuGE workflows. Table 3 summarizes the required information for each component category at each workflow abstraction and each tenet. The fields we select are ultimately an interpretation of our tenet definitions and, as such, are open to adjustment, especially when considering alternate workflow management systems or native software.
Rerun
Rerunning two workflow executions requires matching their logical workflow components. Unsurprisingly, most information captured in this tenet is at the logical layers, only tracking the types and ultimate execution status of each physical component.
Repeat
In addition to the information stored for asserting reruns, repetition stores additional information about component fields (adjustable parameters).
Recompute
Recomputation captures significantly more details of component field values and execution environment. In addition to the component details, we store the logical partitioning of components to machine resources and an execution trace for each processing component.
Reproduce
Reproduction significantly differs, primarily focusing on data components. We store information to assert that data components are configured identically and store the same data summary. How components generate their data summary is a subtlety worth noting; in our case, this is a hash of the data contents but could be implemented as calculating statistical properties or other domain-specific characteristics.
Replicate Scientific/Computational/Total
Replication information is straightforwardly a combination of the Rerun, Recompute, or Repetition information (respectively) and reproduction fields.
Table 3: Tenet data for each DALiuGE component category, workflow abstraction, and reproducibility tenet. In and Out Ports are the DALiuGE abstraction for a dataflow between components. Node and Island define the logical partitioning of components to compute resources to a specific machine and groups of machines, respectively. Status refers to execution status, trace refers to the stack trace of that component, and not all fields are present in all components, as is the case for Filenames in application components, for instance.
Abstraction Tenet Rerun Repeat Recompute Reproduce Replicate-Sci Replicate-Comp Replicate-Total
LGT Field Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control
Type Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y
InPorts Y Y Y Y Y Y Y Y Y - - - Y Y Y Y Y Y Y Y Y
OutPorts Y Y Y Y Y Y Y Y Y - - - Y Y Y Y Y Y Y Y Y
LG Num-CPUs - - - Y - - Y - - - - - - - - Y - - Y - -
Field-Keys - - - Y Y Y Y Y Y - - - - - - Y Y Y Y Y Y
Field-Values - - - - - - Y Y Y - - - - - - Y Y Y - - -
Data-Volume - - - - Y - - Y - - - - - - - - Y - - Y -
Filenames - - - - - - - Y - - - - - - - - Y - - - -
PGS Type Y Y Y Y Y Y Y Y Y - Y - Y Y Y Y Y Y Y Y Y
Name Y - - Y - - Y - - - - - Y - - Y - - Y - -
Storage-Type - Y - - Y - - Y - - Y - - Y - - Y - - Y -
Rank - - - - - - Y Y Y - - - - - - Y Y Y - - -
PGT Node - - - - - - Y Y Y - - - - - - Y Y Y - - -
Island - - - - - - Y Y Y - - - - - - Y Y Y - - -
PG Node-IP-Address - - - - - - Y Y Y - - - - - - Y Y Y - - -
Island-IP-Address - - - - - - Y Y Y - - - - - - Y Y Y - - -
RG Status Y Y - Y Y - Y Y - - Y - Y Y - Y Y - Y Y -
Trace - - - - - - Y - - - - - - - - Y - - - - -
Data-Summary - - - - - - - - - - Y - - Y - - Y - - Y -
5.3 Demonstrative Lowpass Workflow
Figure 2: Generic lowpass filter workflow. The filter component is swapped out for specific implementations, but the rest of the workflow remains the same. Boxes with blue bars represent tasks, the RAM logos represent data storage in memory. The final result is stored in a file.
The workflow used to demonstrate the implemented functionality generates a pure sine wave, injects random Gaussian noise, and then attempts to filter the result with a lowpass filter (based on convolution with a Hann window). While structurally simple, each component is highly configurable and moreover, we implement the core filter component with four different methods; a NumPy [47] pointwise brute-force convolution filter for reference, a NumPy FFT filter, a FFTW [48] FFT filter and PyCuda [49] GPU based FFT filter. Figure 2 contains a single rendering of this workflow. This workflow highlights numerical differences between FFT implementations, a fundamental operation in astronomical sciences and wider scientific computing.
Figure 3: Workflows including the filter task and NCC calculation. Symbols are as in Figure 2. The hard disk icons represent files.
In addition to the basic workflow, we construct an extended workflow that computes the Normalized Cross-Correlation NCC [50] between the original noisy signal and the recovered filtered signal divided by the length of the signal. The resulting metric effectively gives the probability that any filtered sample matches its corresponding sample in the original signal. Figure 3 contains a single rendering of this workflow using the DALiuGE graphical workflow editor. This metric is relative, and therefore, we use the length of the signal to determine the numerical precision required to compare our methods as:
$2^{\lceil log_{2}(L)\rceil}.$ (1)
Our signal length is $512$ in all trials, so we require three digits of precision in our NCC metric. We execute each of the basic and extended workflows 10 times, changing the random seed for the noise generator to the trial number in both cases.
We now outline our example use cases for each tenet:
Rerun
We compare the rerun hashes of each standard workflow. Successful reruns mainly establish trust in the workflow management system itself. While each workflow uses different components, they are all implemented in Python and therefore should be reruns.
Repeat
We compare the repeat hashes of each extended workflow trial. Trials that match can effectively have their results bundled together in concert in making a subsequent claim, such as the efficacy of a filter in this case.
Recompute
We compare the recompute hashes of each workflow trial. Without an identical computing environment, one would not expect any workflow to be a recomputation. Establishing that workflows are not recomputations guarantees differences between them.
Reproduce
For each repeat trial, we compare the data hash of the signal and window data components; the signal components should differ, but the window should remain identical. By comparing the reproduce signature for each trial, we see if each type of filer consistently filters to the same quality and whether this quality differs between each implementation. This test exhibits how building result post-processing into workflows directly allows the precise testing of highly abstract ideas (e.g., the performance of a filter implementation).
Replicate Scientific/Computational/Total
For each type of replication, we compare the hashes of the extended workflow trials. One would expect some repeat trials of the same method to be scientific replicas since the performance of a particular filter should be consistent. A computational replica would be surprising, but establishing total replication on different computing hardware verifies our claims.
6 Demonstration
To demonstrate our methodology as a practical means of comparing workflow designs and implementations, we provide example workflow executions designed to highlight our tenets and real-world examples of where they would apply to astronomical interferometric data processing. All tenet signatures for all referenced workflow trials are available online [51]. Running multiple trials of each workflow, changing only the random noise introduced to the base signal, and comparing our tenet hashes provides a straightforward way to explore what they uncover. Table 4 in appendix A contains the truncated hashes of all plain workflow executions, omitted here for brevity. Table 5 contains the truncated hashes of all NCC test workflow executions, also omitted here for brevity.
For our lowpass filter workflow, we find that although each workflow uses a different library implementation, all trials show the same rerun signature. Repeat trials, differing only in their component parameterization, are identified as such; repeat hashes are the same between implementation methods but differ from their equivalent rerun signatures. In the complete cross-correlation trials, most, but not all, repeat executions are reproductions, indicating that regardless of input data specifics, the filtering component recovers an equivalent amount of signal. The slight variation in signatures indicates that some methods are less effective for some inputs. Upon further inspection, the NCC measure differs by 0.001, which is replicated across all methods on the same input trials. By singling out the input signal data components and comparing their reproduction data hashes, we can confirm that each repeated trial operates on different data. Moreover, despite differing implementations, many cross-correlation repeat trials are reproductions between implementations, demonstrating that each method performs equally well as a filter despite differences in the specific data results. No trials are recomputations, which is to be expected since the runtime trace of each workflow will vary slightly even on the same machine. Scientific replicas single out trials with identical filtering performance and method, formally asserting the equivalence between the FFTW, Cuda FFT, Numpy FFT, and Numpy Pointwise workflows. No trials are computational replicas. Trials matching in workflow and reproductions are total replicas.
6.1 Compute Environment
All trials were executed on a machine with an Intel Core 17-10875H, Nvidia RTX 2080 Super (Mobile), and 64GB of RAM. Experiments were run with DALiuGE version 2.3.2, EAGLE version 4.5.2, and lowpass components version 0.2.
7 Conclusion
Computational workflows power contemporary data-intensive sciences. Workflows play an increasingly important role in astronomical sciences in particular. To build scientific trust in these complex software systems, research engineers adopt industry-standard software development methodologies to implement and test the software. We provide an additional methodology for testing and verifying consistency between and across execution runs using either the same or varying implementations of specific algorithms. We construct a hash-graph-based signature mechanism termed BlockDAGs to characterize workflow executions with respect to these standards and add the resulting hash keys to the provenance information of individual runs. Our demonstrative lowpass filter example shows that this method asserts the scientific equivalence of four different implementations of the same basic workflow while simultaneously identifying the output-data differences between them. Workflow management systems guide workflow execution and design, and our method to test the scientific equality between workflow executions allows us to verify that these systems deliver on their promised benefits. Additional considerations, such as adding formal provenance data to the component hashes and assessing the impact of hardware failure on large scientific workflow instances, would further verify the efficacy of our method. Implementing our BlockDAG mechanism in other workflow management systems would allow for workflow verification between workflow management systems. By providing a test-driven method to determine the scientific equivalence between workflow executions we provide verifiable evidence that next-generation science instruments, such as the SKA produces high quality, reproducible results.
8 Software and data availability
All components used in this study are publicly available as a Python package [52]. The workflows themselves [53], and all data products from each workflow trial execution are available online [51]. The DALiuGE system [54] and the DALiuGE graph editor EAGLE [55] are open source and available as well.
CRediT authorship contribution statement
Nicholas J. Pritchard: Conceptualization, Methodology, Software, Validation, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing. Andreas Wicenec: Conceptualization, Resources, Writing - Review & Editing, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgements
This research was supported by an Australian Government Research Training Program (RTP) Scholarship.
Appendix A Workflow trial signatures
Here, we present truncated signatures output from all trial runs referred to in section 6. Table 4 contains the signature values for plain lowpass filter tests. Table 5 contains the signatures from the extended NCC trials.
Table 4: Table of plain workflow signatures. Truncated from the original signature values to the first five characters for brevity. Signature standards are labeled as follows: RR - Rerun, RT - Repeat, RC - Recompute, RP - Reproduce, RPL-S - Replicate Scientifically, RPL-C - Replicate Computationally, RPL-T - Replicate Totally. Workflow trial names formatted as method-library-trial.
Workflow Trial RR RT RC RP RPL-S RPL-C RPL-T
FFT-fftw-0 3cabf deb85 d5556 4fa43 f9615 f321d ca359
FFT-fftw-1 3cabf deb85 117c4 9ae2d bfc4b 1e3cf 4d937
FFT-fftw-2 3cabf deb85 03d3c 26e2a 8b785 17ca2 7fb60
FFT-fftw-3 3cabf deb85 3bc5e 0d4a1 6648d d0034 687bf
FFT-fftw-4 3cabf deb85 58e17 13c75 33366 c433b 6d1fe
FFT-fftw-5 3cabf deb85 e1a4b 2d0d5 4ce42 56002 b4138
FFT-fftw-6 3cabf deb85 284d6 b345f 8c086 fa69b 589bb
FFT-fftw-7 3cabf deb85 a5f11 dd944 5959c 882e0 58b64
FFT-fftw-8 3cabf deb85 c8e02 7b758 37c0d 3f3c5 07808
FFT-fftw-9 3cabf deb85 d9d6b 0da65 7c3fb 29eb0 b6bcc
Pointwise-np-0 3cabf 2f7f9 7cb68 3449a 837f2 70370 7549c
Pointwise-np-1 3cabf 2f7f9 080d7 928a7 aa7a0 45b3a 5b6a9
Pointwise-np-2 3cabf 2f7f9 24963 59f82 305bd 36c5f c7ce8
Pointwise-np-3 3cabf 2f7f9 6fa7a efb06 c2896 ef68f 2e278
Pointwise-np-4 3cabf 2f7f9 4b180 b3742 c0e64 5b7bb 4eb40
Pointwise-np-5 3cabf 2f7f9 3e519 f07ec 65a85 0a35f 00308
Pointwise-np-6 3cabf 2f7f9 0c776 501eb b3fae 8d837 3d6b0
Pointwise-np-7 3cabf 2f7f9 6d9bc 6ebb8 d4f24 60547 f4bd6
Pointwise-np-8 3cabf 2f7f9 4019e b9cf6 fb869 2c487 d09bc
Pointwise-np-9 3cabf 2f7f9 91d11 5ce91 3be4d b2ae1 9c36d
FFT-np-0 8619d 6bc99 76aac 8fd08 faf14 eac3f 07d3a
FFT-np-1 8619d 6bc99 326f1 bf29b 52ae2 3748e 9b208
FFT-np-2 8619d 6bc99 c493b bf29b 52ae2 64fa3 9b208
FFT-np-3 8619d 6bc99 09956 8fd08 faf14 dc375 07d3a
FFT-np-4 8619d 6bc99 e15ba bf29b 52ae2 2a5df 9b208
FFT-np-5 8619d 6bc99 c261c bf29b 52ae2 ef2c5 9b208
FFT-np-6 8619d 6bc99 a3cc5 bf29b 52ae2 a59c6 9b208
FFT-np-7 8619d 6bc99 7295c bf29b 52ae2 72836 9b208
FFT-np-8 8619d 6bc99 a03e0 8fd08 faf14 3131e 07d3a
FFT-np-9 8619d 6bc99 88ff5 bf29b 52ae2 6d378 9b208
FFT-cuda-0 8619d 85f98 09b82 8fd08 faf14 9343a 7ecb0
FFT-cuda-1 8619d 85f98 0ebd4 bf29b 52ae2 2df4d 6f8a9
FFT-cuda-2 8619d 85f98 641dc bf29b 52ae2 78071 6f8a9
FFT-cuda-3 8619d 85f98 7e4bf 8fd08 faf14 d7105 7ecb0
FFT-cuda-4 8619d 85f98 59a07 bf29b 52ae2 4f18d 6f8a9
FFT-cuda-5 8619d 85f98 fb74a bf29b 52ae2 46996 6f8a9
FFT-cuda-6 8619d 85f98 6b8ea bf29b 52ae2 25474 6f8a9
FFT-cuda-7 8619d 85f98 8581f bf29b 52ae2 a9abf 6f8a9
FFT-cuda-8 8619d 85f98 900e1 bf29b 52ae2 93555 6f8a9
FFT-cuda-9 8619d 85f98 024b7 bf29b 52ae2 1dfe3 6f8a9
Table 5: Table of NCC workflow signatures. Truncated from the original signature values to the first five characters for brevity. Signature standards are labeled as follows: RR - Rerun, RT - Repeat, RC - Recompute, RP - Reproduce, RPL-S - Replicate Scientifically, RPL-C - Replicate Computationally, RPL-T - Replicate Totally. Workflow trial names formatted as method-library-trial.
Workflow Trial RR RT RC RP RPL-S RPL-C RPL-T
Pointwise-np-ncc-0 3cabf deb85 d5556 4fa43 f9615 f321d ca359
Pointwise-np-ncc-1 3cabf deb85 117c4 9ae2d bfc4b 1e3cf 4d937
Pointwise-np-ncc-2 3cabf deb85 03d3c 26e2a 8b785 17ca2 7fb60
Pointwise-np-ncc-3 3cabf deb85 3bc5e 0d4a1 6648d d0034 687bf
Pointwise-np-ncc-4 3cabf deb85 58e17 13c75 33366 c433b 6d1fe
Pointwise-np-ncc-5 3cabf deb85 e1a4b 2d0d5 4ce42 56002 b4138
Pointwise-np-ncc-6 3cabf deb85 284d6 b345f 8c086 fa69b 589bb
Pointwise-np-ncc-7 3cabf deb85 a5f11 dd944 5959c 882e0 58b64
Pointwise-np-ncc-8 3cabf deb85 c8e02 7b758 37c0d 3f3c5 07808
Pointwise-np-ncc-9 3cabf deb85 d9d6b 0da65 7c3fb 29eb0 b6bcc
FFT-fftw-ncc-0 3cabf 2f7f9 7cb68 3449a 837f2 70370 7549c
FFT-fftw-ncc-1 3cabf 2f7f9 080d7 928a7 aa7a0 45b3a 5b6a9
FFT-fftw-ncc-2 3cabf 2f7f9 24963 59f82 305bd 36c5f c7ce8
FFT-fftw-ncc-3 3cabf 2f7f9 6fa7a efb06 c2896 ef68f 2e278
FFT-fftw-ncc-4 3cabf 2f7f9 4b180 b3742 c0e64 5b7bb 4eb40
FFT-fftw-ncc-5 3cabf 2f7f9 3e519 f07ec 65a85 0a35f 00308
FFT-fftw-ncc-6 3cabf 2f7f9 0c776 501eb b3fae 8d837 3d6b0
FFT-fftw-ncc-7 3cabf 2f7f9 6d9bc 6ebb8 d4f24 60547 f4bd6
FFT-fftw-ncc-8 3cabf 2f7f9 4019e b9cf6 fb869 2c487 d09bc
FFT-fftw-ncc-9 3cabf 2f7f9 91d11 5ce91 3be4d b2ae1 9c36d
FFT-np-ncc-0 8619d b09c7 a6758 8fd08 faf14 c6cb6 89a7c
FFT-np-ncc-1 8619d b09c7 d5ed5 bf29b 52ae2 b21bb cfb27
FFT-np-ncc-2 8619d b09c7 20176 bf29b 52ae2 84b38 cfb27
FFT-np-ncc-3 8619d b09c7 ecc8b 8fd08 faf14 585e2 89a7c
FFT-np-ncc-4 8619d b09c7 4a455 bf29b 52ae2 17d62 cfb27
FFT-np-ncc-5 8619d b09c7 10797 bf29b 52ae2 3119d cfb27
FFT-np-ncc-6 8619d b09c7 147db bf29b 52ae2 4ae6b cfb27
FFT-np-ncc-7 8619d b09c7 d181b bf29b 52ae2 1962b cfb27
FFT-np-ncc-8 8619d b09c7 3bc4c 8fd08 faf14 2f68d 89a7c
FFT-np-ncc-9 8619d b09c7 e9262 bf29b 52ae2 9518f cfb27
FFT-cuda-ncc-0 8619d 85f98 09b82 8fd08 faf14 9343a 7ecb0
FFT-cuda-ncc-1 8619d 85f98 0ebd4 bf29b 52ae2 2df4d 6f8a9
FFT-cuda-ncc-2 8619d 85f98 641dc bf29b 52ae2 78071 6f8a9
FFT-cuda-ncc-3 8619d 85f98 7e4bf 8fd08 faf14 d7105 7ecb0
FFT-cuda-ncc-4 8619d 85f98 59a07 bf29b 52ae2 4f18d 6f8a9
FFT-cuda-ncc-5 8619d 85f98 fb74a bf29b 52ae2 46996 6f8a9
FFT-cuda-ncc-6 8619d 85f98 6b8ea bf29b 52ae2 25474 6f8a9
FFT-cuda-ncc-7 8619d 85f98 8581f bf29b 52ae2 a9abf 6f8a9
FFT-cuda-ncc-8 8619d 85f98 900e1 bf29b 52ae2 93555 6f8a9
FFT-cuda-ncc-9 8619d 85f98 024b7 bf29b 52ae2 1dfe3 6f8a9